OpenStudy (anonymous):

If a*b= a^2-b^2/a+b, what is 3*2.

OpenStudy (anonymous):

-5 -2 1 2 5

OpenStudy (anonymous):

I got 1

OpenStudy (jamesj):

So this expression is defining a new operation: * a*b is defined to be (a^2 - b^2)/(a+b) is that the right placement of brackets?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Is my answer 1, right?

OpenStudy (jamesj):

Well, notice that \[ a^2 - b^2 = (a-b)(a+b) \] Hence \[ \frac{a^2-b^2}{a+b} = \frac{(a+b)(a-b)}{a+b} = a-b \] Thus 3*2 = 3 - 2 = 1, which is what you wrote above.

OpenStudy (anonymous):

So when you get problems like this, you are basically plugging in the values into where the c and f notations?

OpenStudy (anonymous):

like you plug in the numbers where they are supposed to go? its that easy?

OpenStudy (jamesj):

Yes. But it's worth looking at the expression to see if you can simplify it first, as I did above.

OpenStudy (anonymous):

Do u mind looking at this one?

OpenStudy (jamesj):

post as a new question if you don't mind.

OpenStudy (anonymous):

Let r (triangle) s=rs + s for all integers r and s. What is the value of 4 (triangle) 5?

OpenStudy (anonymous):

@JamesJ

OpenStudy (anonymous):

options: 18 20 24 25 40

OpenStudy (jamesj):

what do you think it is?

OpenStudy (anonymous):

I got 25:)?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

I'll post more if I have to in a sec!